# Depth-Breadth Synergy in RLVR: Unlocking LLM Reasoning Gains with Adaptive Exploration

### v4

@@ Abstract
Reinforcement Learning with Verifiable Reward (RLVR) has emerged as a powerful paradigm for unlocking reasoning capabilities in large language models, yet its full potential is hindered by two under-explored dimensions: Depth—the hardest problem a model can sample; Breadth—the number of instances consumed in a single iteration. We dissect the popular GRPO algorithm and reveal a systematic bias: the cumulative-advantage disproportionately weights samples with medium accuracy, while down-weighting the low-accuracy instances that are crucial for pushing reasoning boundaries. To rectify the depth neglect, we introduce Difficulty Adaptive Rollout Sampling (DARS), which re-weights hard problems through targeted multi-stage rollouts, thereby increasing the number of positive rollouts for hard problems. Empirically, naively enlarging rollout size only accelerates convergence and even hurts Pass@K. Our DARS, in contrast, delivers consistent Pass@K gains without extra inference cost at convergence. Just as we adaptively expanded the depth of exploration, we now ask whether aggressively scaling the breadth of training data can further amplify reasoning gains. To this end, we intensely scale batch size and replace PPO’s mini-batch iterations with full-batch updates over multiple epochs. Increasing breadth significantly enhances Pass@1 performance. Large-breadth training sustains high token-level entropy, indicating continued exploration and reduced gradient noise. We further present DARS-Breadth, which augments DARS with large breadth, and demonstrate simultaneous gains in Pass@K and Pass@1. The results confirm that breadth and adaptive exploration across depth operate as orthogonal dimensions in RLVR, which are key to unleashing the reasoning power of RLVR.
Figure 1: Training dynamics of Pass@1 and Pass@K performance. We show that our DARS significantly improves Pass@K performance and is complementary to breadth scaling to further improve Pass@1 performance.
@@ 1 Introduction
The emergence of reasoning-centric Large Language Models (LLMs) exemplified by OpenAI-o1 (Jaech et al., 2024), DeepSeek-R1 (Guo et al., 2025), and Kimi-1.5 (Team et al., 2025), has pushed the frontier of LLM capability, especially for demanding tasks in complex reasoning such as mathematics and programming. Unlike conventional instruction tuning that relies on human-labeled data or RLHF pipelines that demand an auxiliary, well-trained reward model (Ouyang et al., 2022a; Achiam et al., 2023; Grattafiori et al., 2024), this leap is driven by large-scale Reinforcement Learning with Verifiable Rewards (RLVR; Guo et al. 2025; Zeng et al. 2025) for which correctness can be automatically and deterministically checked. The rewards of RLVR are granted solely when a model’s output matches the ground-truth mathematical answer or passes all unit tests for code, allowing scalable verification without manual labeling. RLVR is now regarded as a promising path toward self-evolving LLMs, potentially bringing us closer to more powerful intelligence.
@@
However, existing RLVR frameworks inadequately address the interplay between exploration depth (difficulty scaling) and breadth (iteration instance quantity scaling), which leads to insufficient performance gain for both Pass@1 and Pass@K. In this paper, we conduct a systematic analysis of two under-exploited dimensions in RLVR: Depth and Breadth.
@@
For the dimension of depth, our investigation reveals that existing methods of GRPO (Shao et al., 2024) and its variants (Yu et al., 2025; Liu et al., 2025b), while adept at estimating the advantage of a single rollout, are undermined by a distorted cumulative advantage at the group level. This distortion disproportionately allocates attention to instances of medium difficulty, neglecting high-difficulty instances indispensable for complex reasoning, as illustrated in Figure 3. This bias fundamentally limits depth, the hardest problems a model can learn to solve, and constrains Pass@K performance. To counteract this depth neglect, we propose Difficulty-Adaptive Rollout Sampling (DARS). DARS performs a lightweight first-stage rollout to estimate per-problem accuracies, then allocates additional compute via targeted multi-stage rollouts to low-accuracy problems. By expanded sampling on hard problems, DARS re-weights the cumulative advantage, making it easier for LLMs to learn ‘deep’ samples and improving Pass@K performance.
@@
We further identify breadth as the instance quantity consumed in a single iteration. We observe that breadth has a significant impact on the LLM’s performance and continuous exploration capability, as shown in Figure 5. We significantly increase the training batch size and replace PPO-minibatch updates with full-batch updates for multiple PPO-epochs. This seemingly simple change dramatically improves Pass@1 and sustains high token-level entropy throughout training, suggesting that breadth acts as implicit entropy regularization that delays premature convergence. Importantly, the gains from breadth are complementary to those from depth: we present DARS-Breadth that combines our DARS with large-breadth training, producing simultaneous boosts in both Pass@K and Pass@1.
Our contributions can be summarized as follows:
@@
  • We conduct a systematic analysis on depth and breadth in RLVR, and uncover the depth bias in GRPO: cumulative advantage silently underweights low-accuracy, high-difficulty samples, capping Pass@K performance.
@@ 2 Understanding RLVR from Depth and Breadth
In this section, we identify the two dimensions of Depth and Breadth and examine their impact on the Pass@1 and Pass@K metrics during the RLVR process. For training details and specific experimental settings, please refer to Section 4.1.
@@ 2.1 Depth: The Hardest Problem Sampled in RLVR
We first identify Depth as the hardest problem that can be correctly answered in the RLVR training process. In the GRPO training process, groups whose entire rollouts yield incorrect answers suffer from gradient vanishing. Hence, sampling high-difficulty questions with correct reasoning paths is crucial for LLM training. In this subsection, we first show that merely increasing rollout size does not consistently yield significant gains in Pass@K performance, and sometimes can even be harmful. We then quantify GRPO’s cumulative advantage and highlight its under-weighting of high-difficulty samples.
Figure 2: Training dynamics of Pass@1 and Pass@K performance of Qwen2.5-Math-1.5b and Qwen2.5-Math-7b with different rollout size.
Naive Scaling of Rollout Size Benefits Pass@1, But Not Necessarily Pass@K.
We present the training dynamics of Pass@1 and Pass@K performance during the RLVR training process in Figure 2. Enlarging the rollout size allows the sampling of correct solutions to hard problems during training. We originally assumed this would benefit Pass@K performance; however, experimental results show that this is not always the case. We find that Qwen2.5-Math-7b can significantly benefit from an increased rollout size, whereas for Qwen2.5-Math-1.5b, naively scaling rollout size can even harm Pass@K performance.
Cumulative Advantage Bias in GRPO Variants hinders the improvement of Pass@K.
In the GRPO framework, the advantage estimation is derived by normalizing binary rewards:
@@
where $r_{i}$ is the binary reward of $i_{th}$ rollout, $u$ is the mean value of the group rewards $u=\text{mean}(\{R_{i}\}_{i=1}^{G})$ and $\sigma$ is the standard deviation of the group rewards $\sigma=\text{std}(\{R_{i}\}_{i=1}^{G})$. In the case of binary rewards, $u$ also represents the accuracy of LLM rollouts. Dr. GRPO  (Liu et al., 2025b) removes the standard-deviation term from the advantage computation to eliminate question-level difficulty bias, and demonstrates its superiority through extensive experiments. Consequently, the experiments reported in this study were conducted primarily though the Dr. GRPO methodology. We show more results of std-based advantage in Appendix C.1.
@@
$\mathcal{A}_{\text{group}}=2Nu(1-u),$ (2)
@@ 2.2 Breadth: Iteration Instance Quantity in RLVR
We identify Breadth as the quantity of instances used in each iteration of the RLVR process. In this subsection, we discuss the Pass@1 performance benefits of simply increasing the batch size for the RLVR process. In section 4.3, we further discussed the impact of replacing the PPO mini batch with the full batch update for multiple PPO epochs.
Breadth Matters for Pass@1 Performance.
Most studies (Liu et al., 2025b; a; Yan et al., 2025; Fu et al., 2025) conventionally set the batch size to 128. In this subsection, we drastically increase the batch size to 3072 and plot the training dynamics of Pass@1 and Pass@32 performance in Figure 4. Naively increasing the batch size brings a Pass@1 improvement for all models, yet it harms the Pass@128 performance of Qwen2.5-Math-1.5b. We consider that increasing the quantity of instances used in each iteration makes the gradient direction more accurate and reduces the impact of noise, thereby improving Pass@1 performance.
Figure 4: Training dynamics of Pass@1 and Pass@K performance of Qwen2.5-Math-1.5b and Qwen2.5-Math-7b with different batch size.
@@
We further analyze the relationship between Pass@1 and token entropy during the training process, as shown in Figure 5. Higher token entropy signifies stronger exploration capabilities in LLMs. Increasing training breadth allows an LLM to achieve higher entropy at the same Pass@1 accuracy. We consider that large breadth serves as an entropy regularization mechanism, preventing premature convergence and boosting Pass@1 while sustaining entropy.
@@ 3.1 Difficulty Adaptive Rollout Sampling (DARS)
$\mathcal{A}_{\text{group}}^{HW}(q_{j})=2(1-\hat{a}_{j})\mathcal{A}_{\text{group}}^{N^{pre}}(0.5).$ (7)
@@
$\Delta n_{j}^{\text{HW}}=\mathrm{min}(\left\lceil\frac{2x_{j}\cdot\mathcal{A}_{\text{group}}^{N^{pre}}(0.5)-\mathcal{A}_{\text{group}}^{N^{pre}}(\hat{a}_{j})}{\mathcal{S}(\hat{a}_{j})}\right\rceil,N^{\mathrm{max}}-N^{pre}).$ (8)
@@ 3.2 Depth Synergy with Breadth Scaling
3.3 Training Target
$\mathcal{J}(\theta)=$ $\mathbb{E}_{(q\sim\mathcal{D},\{o_{i}\}_{i=1}^{\mathcal{G}}\sim\pi_{\theta_{\text{old}}}(q)}$ (9)
$\Bigg[\frac{1}{G}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\Bigg(\min\Big(r_{i,t}(\theta)\hat{A}_{i,t},\ \text{clip}\Big(r_{i,t}(\theta),1-\varepsilon,1+\varepsilon\Big)\hat{A}_{i,t}\Big))\Bigg)\Bigg],$
where
The token advantage $\hat{A}_{i,t}$ is computed using Equation 1.
4 Experiments
4.1 Setup
Evaluation and Training.
We evaluate the RLVR process using 5 widely used mathematical reasoning benchmarks: MATH-500 (Lightman et al., 2023), OlympiadBench (He et al., 2024), MinvervaMath (Lewkowycz et al., 2022), AIME24, and AMC23. We combine all of the evaluation benchmarks to report Pass@1 (Avg@128) and Pass@K performance. The training data used in this work is OpenR1-45K, which is a subset of OpenR1-Math-220k (Hugging Face, 2025). More details are shown in Appendix B.1.
Baseline and Methods:
We compare the following methods: (1) Other Post-Training Models: The Qwen2.5-Math-Instruct (Yang et al., 2024) series, SimpleRL-Zero-7B Zeng et al. (2025) model, and Oat-Zero (Liu et al., 2025b) series. (2) RLVR-baseline: Dr. GRPO with rollout size 8 and batch size 128. (3) Depth-Naive: Simply increasing the rollout size to 32. (4) Breadth-Naive: Simply increasing the batch size to 3072. (5) DARS-ET/HW: Our algorithm introduced in Section 3.1 with Equal-Treat/Hardness-Weighted schedule, using batch size 128 and $N^{\max}=32$. (6) DARS-ET/HW-Breadth: Our Depth-and-Breadth synergy algorithm introduced in Section 3.2, using batch size 3072 and $N^{\max}=32$. For all methods, the number of PPO mini-steps is uniformly set to 2.
Evaluation Protocol:
For all baseline methods, we select the checkpoint with the best Pass@1 performance for reporting. For our DARS method, we selected the checkpoint that achieved the best Pass@128 performance after surpassing the baseline Pass@1 performance. Table 1 summarizes the Avg@128 performance on each benchmark, the overall Pass@1 across all test data, and the Pass@128 performance.
@@ 4.1 Setup
Qwen2.5-Math-1.5B as the Base Model
Qwen2.5-Math-1.5B-Base 4.0 35.1 16.2 20.8 9.5 21.1 77.9
Qwen2.5-Math-1.5B-Ins 10.2 67.9 34.7 42.6 24.4 43.5 79.9
Oat-Zero-1.5B 16.4 73.0 35.5 47.4 26.8 46.3 80.3
Our RLVR Replication 14.7 75.9 39.4 47.5 31.2 49.6 79.6
@@
Qwen2.5-Math-7B as the Base Model
Qwen2.5-Math-7B-Base 11.6 52.3 19.7 35.2 15.3 30.1 82.1
Qwen2.5-Math-7B-Ins 12.9 81.5 39.9 47.0 34.1 52.0 82.3
SimpleRL-Zero-7B 23.3 72.8 36.1 52.8 26.8 46.9 82.5
Oat-Zero-7B 31.3 79.2 42.5 59.4 33.7 53.4 79.7
Our RLVR Replication 26.8 82.2 44.3 57.2 35.7 55.3 81.4

### v5

@@ Abstract
Reinforcement Learning with Verifiable Reward (RLVR) is a powerful method for enhancing the reasoning abilities of Large Language Models, but its full potential is limited by a lack of exploration in two key areas: Depth (the difficulty of problems) and Breadth (the number of training instances). Our analysis of the popular GRPO algorithm reveals a bias that down-weights difficult, low-accuracy problems, which are crucial for improving reasoning skills. To address this, we introduce Difficulty Adaptive Rollout Sampling (DARS), a method that re-weights difficult problems by using targeted, multi-stage rollouts. DARS increases the number of rollout outcomes for these harder problems according to our proposed re-balancing schedules and leads to consistent gains in Pass@K. We discovered that increasing rollout size alone does not improve performance and may actually impair it. In contrast, scaling the batch size to increase breadth via full-batch updates significantly boosted Pass@1 metrics. This improvement stems from higher token-level entropy, ensuring robust exploration and minimized gradient noise. Finally, we present DARS-Breadth, a combined approach that uses DARS with a large breadth of training data. This method demonstrates simultaneous gains in both Pass@K and Pass@1, confirming that depth (adaptive exploration) and breadth (scaling the training data) are orthogonal and essential dimensions for unlocking the full power of RLVR.
Machine Learning, ICML
yangzhch6@gmail
Project Repo: https://github.com/yangzhch6/DARS
@@ 1 Introduction
The emergence of reasoning-centric Large Language Models (LLMs) exemplified by OpenAI-o1 (Jaech et al., 2024), DeepSeek-R1 (Guo et al., 2025), and Kimi-1.5 (Team et al., 2025), has pushed the frontier of LLM capability, especially for demanding tasks in complex reasoning such as mathematics and programming. Unlike conventional instruction tuning that relies on human-labeled data or RLHF pipelines that demand an auxiliary, well-trained reward model (Ouyang et al., 2022b; Achiam et al., 2023; Grattafiori et al., 2024), this leap is driven by large-scale Reinforcement Learning with Verifiable Rewards (RLVR; Guo et al. 2025; Zeng et al. 2025) for which correctness can be automatically and deterministically checked. The rewards of RLVR are granted solely when a model’s output matches the ground-truth mathematical answer or passes all unit tests for code, allowing scalable verification without manual labeling. RLVR is now regarded as a promising path toward self-evolving LLMs, potentially bringing us closer to more powerful intelligence.
@@
However, existing RLVR frameworks inadequately address the interplay between exploration depth (difficulty scaling) and breadth (iteration instance quantity scaling), which leads to insufficient performance gain for both Pass@1 and Pass@K. In this paper, we conduct a systematic analysis of these two under-exploited dimensions in RLVR: Depth and Breadth.
@@
For the dimension of depth, our investigation reveals that existing methods of GRPO (Shao et al., 2024) and its variants (Yu et al., 2025; Liu et al., 2025b), while adept at estimating the advantage of a single rollout, are undermined by a distorted cumulative advantage at the group level. This distortion disproportionately allocates attention to instances of medium difficulty, neglecting high-difficulty instances indispensable for complex reasoning, as illustrated in Figure 1. This bias fundamentally limits depth, the hardest problems a model can learn to solve, and constrains Pass@K performance. To counteract this depth neglect, we propose Difficulty-Adaptive Rollout Sampling (DARS). DARS performs a lightweight first-stage rollout to estimate per-problem accuracies, then allocates additional compute via targeted multi-stage rollouts to low-accuracy problems. By expanding sampling on hard problems, DARS re-weights the cumulative advantage, making it easier for LLMs to learn ‘deep’ samples and improving Pass@K performance.
@@
We further identify breadth as the instance quantity consumed in a single iteration. We observe that breadth has a significant impact on the LLM’s performance and continuous exploration capability, as shown in Figure 4. We significantly increase the training batch size and replace PPO-minibatch updates with full-batch updates for multiple PPO-epochs. This seemingly simple change dramatically improves Pass@1 and sustains high token-level entropy throughout training, suggesting that breadth acts as implicit entropy regularization that delays premature convergence. Importantly, the gains from breadth are complementary to those from depth: we present DARS-Breadth that combines our DARS with large-breadth training, producing simultaneous boosts in both Pass@K and Pass@1. Our contributions can be summarized as follows:
@@
  • We conduct a systematic analysis on depth and breadth in RLVR, and uncover the depth bias in GRPO: cumulative advantage underweights low-accuracy, high-difficulty samples, capping Pass@K performance.
@@ 2 Understanding RLVR from Depth and Breadth
Figure 2: Training dynamics of Pass@1 and Pass@K performance of Qwen2.5-Math-1.5B and -7B with different rollout size.
@@ 2.1 Depth: The Hardest Problem Sampled in RLVR
We first identify Depth as the hardest problem that can be correctly answered in the RLVR training process. In the GRPO training process, groups whose entire rollouts yield incorrect answers suffer from gradient vanishing. Hence, sampling high-difficulty questions with correct reasoning paths is crucial for LLM training. We first show that merely increasing rollout size does not consistently yield significant gains in Pass@K performance, and sometimes can even be harmful. We then quantify GRPO’s cumulative advantage and highlight its under-weighting of high-difficulty samples.
Naive Scaling of Rollout Size Benefits Pass@1, But Not Necessarily Pass@K. We present the training dynamics of Pass@1 and Pass@K performance during the RLVR training process in Figure 2. Enlarging the rollout size allows the sampling of correct solutions to hard problems during training. We originally assumed this would benefit Pass@K performance; however, experimental results show that this is not always the case. We find that Qwen2.5-Math-7B can significantly benefit from an increased rollout size, whereas for Qwen2.5-Math-1.5b, naively scaling rollout size can even harm Pass@K performance.
Cumulative Advantage Bias in GRPO Variants hinders the improvement of Pass@K. In the GRPO framework, the advantage estimation is derived by normalizing binary rewards:
@@
where $r_{i}$ is the binary reward of $i_{th}$ rollout, $u$ is the mean value of the group rewards $u=\text{mean}(\{R_{i}\}_{i=1}^{G})$ and $\sigma$ is the standard deviation of the group rewards $\sigma=\text{std}(\{R_{i}\}_{i=1}^{G})$. In the case of binary rewards, $u$ also represents the accuracy of LLM rollouts. Dr. GRPO  (Liu et al., 2025b) removes the standard-deviation term from the advantage computation to eliminate question-level difficulty bias, and demonstrates its superiority through extensive experiments. Consequently, the experiments reported in this study were conducted primarily though the Dr. GRPO methodology. We show more results of std-based advantage in Appendix D.
Figure 3: Training dynamics of Pass@1 and Pass@K performance of Qwen2.5-Math-1.5B and -7B with different batch size.
@@
$\mathcal{A}_{\text{group}}=2Nu(1-u).$ (2)
@@ 2.2 Breadth: Iteration Instance Quantity in RLVR
We define Breadth as the number of instances used per iteration of the RLVR process. We’ll show how increasing the batch size for the RLVR process improves the Pass@1 performance.
Breadth Matters for Pass@1 Performance. Most studies (Liu et al., 2025b, a; Yan et al., 2025; Fu et al., 2025) conventionally set the batch size to 128. In this subsection, we drastically increase the batch size to 3072 and plot the training dynamics of Pass@1 and Pass@32 performance in Figure 3. Naively increasing the batch size brings a Pass@1 improvement for all models, yet it harms the Pass@128 performance of Qwen2.5-Math-1.5B. We consider that increasing the quantity of instances used in each iteration makes the gradient direction more accurate and reduces the impact of noise, thereby improving Pass@1 performance.
@@
High token entropy in LLMs indicates strong exploration capabilities. Our analysis shows a relationship between Pass@1 and token entropy during training. As illustrated in Figure 4, increased training breadth enables LLMs to achieve higher entropy at a given Pass@1 accuracy. We believe a large training breadth acts as a form of entropy regularization, preventing premature convergence and boosting Pass@1 performance while maintaining high entropy.
@@ 3.1 Difficulty Adaptive Rollout Sampling (DARS)
$\mathcal{A}_{\text{group}}^{HW}(q_{j})=2(1-x_{j})\mathcal{A}_{\text{group}}^{N^{pre}}(0.5).$ (7)
@@
$\Delta n_{j}^{\text{HW}}=$ $\mathrm{min}(\left\lceil\frac{2x_{j}\cdot\mathcal{A}_{\text{group}}^{N^{pre}}(0.5)-\mathcal{A}_{\text{group}}^{N^{pre}}(\hat{a}_{j})}{\mathcal{S}(\hat{a}_{j})}\right\rceil,$ (8)
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~N^{\mathrm{max}}-N^{pre}).$
@@ 3.2 Depth Synergy with Breadth Scaling
Qwen2.5-Math-1.5B 4.0 35.1 16.2 20.8 9.5 21.1 77.9
RLVR baseline 14.7 75.9 39.4 47.5 31.2 49.6 79.6
@@
Qwen2.5-Math-7B 11.6 52.3 19.7 35.2 15.3 30.1 82.1
RLVR baseline 26.8 82.2 44.3 57.2 35.7 55.3 81.4
@@
3.3 Training Target
$\mathcal{J}(\theta)=\mathbb{E}_{(q\sim\mathcal{D},\{o_{i}\}_{i=1}^{\mathcal{G}}\sim\pi_{\theta_{\text{old}}}(q)}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\Bigg($ (9)
$~~~~~\min\Big(r_{i,t}(\theta)\hat{A}_{i,t},\ \text{clip}\Big(r_{i,t}(\theta),1-\varepsilon,1+\varepsilon\Big)\hat{A}_{i,t}\Big)\Bigg)\Bigg],$
where
The token advantage $\hat{A}_{i,t}$ is computed using Equation 1.
4 Experiments
4.1 Setup
Evaluation and Training: We evaluate the RLVR process using 5 widely used mathematical reasoning benchmarks: MATH-500 (Lightman et al., 2023), OlympiadBench (He et al., 2024), MinvervaMath (Lewkowycz et al., 2022), AIME24, and AMC23. We combine all of the evaluation benchmarks to report Pass@1 (Avg@128) and Pass@K performance. The training data used in this work is OpenR1-45K, which is a subset of OpenR1-Math-220k (Hugging Face, 2025). Moreover, we adopt the same unbiased, low-variance estimator for Pass@K as used in prior works (Yue et al., 2025a; Chen et al., 2021):
$\text{pass@}K=\mathbb{E}_{x_{i}\sim\mathcal{D}}\left[1-\frac{\binom{n-c_{i}}{k}}{\binom{n}{k}}\right$].
Specifically, when $K=N$, the metric become: $\text{pass@}K=c_{1}\lor c_{2}\lor\cdots c_{128}$.
Baseline and Methods: We compare with: (1) RLVR-baseline: Dr. GRPO with rollout size 8 and batch size 128. (2) Depth-Naive: Simply increasing the rollout size to 32. (3) Breadth-Naive: Simply increasing the batch size to 3072. (4) DARS-ET/HW: Our algorithm introduced in Section 3.1 with Equal-Treat/Hardness-Weighted schedule, using batch size 128 and $N^{\max}=32$. (5) DARS-ET/HW-Breadth: Our Depth-and-Breadth synergy algorithm introduced in Section 3.2, using batch size 3072 and $N^{\max}=32$. For all methods, the number of PPO mini-steps is uniformly set to 2.
Evaluation Protocol: For all baselines, we select the checkpoint with the best Pass@1 performance for reporting. For DARS, we selected the checkpoint that achieved the best Pass@128 performance after surpassing the baseline Pass@1 performance. Table 1 summarizes the Avg@128 performance on each benchmark, the overall Pass@1 across all test data, and the Pass@128 performance.
Figure 7: Training dynamics of Pass@32/Pass@128 and Pass@1 with different training steps of Qwen2.5-Math-1.5B and -7B.